OpenStudy (pphalke):

Find the total monthly payment for a $205,000 mortgage loan at 5.25% for 20 years. The assessed value of the home is $240,000. The annual taxes on the home are 1.25% of the assessed value, and the insurance on the home costs $450 per year I have tried with the calculator its not helping please assist me I will fan medal

OpenStudy (wolf1728):

Do you have an answer for this? (I'm guessing you have some idea of what the answer should be)

OpenStudy (pphalke):

I have no idea but I think its $1668.88 I don't know how I got it

OpenStudy (wolf1728):

We will first have to calculate the insurance and tax and add that into the mortgage total.

OpenStudy (pphalke):

how do I go about doing that

OpenStudy (wolf1728):

The assessed value is 240,000 and the annual taxes are 1.25% for 20 years that would be .0125 * 240,000 * 20 = $60,000 20 years of insurance = 450 * 20 = $9,000 So, I think the entire mortgage should be for 205,000 + 60,000 + 9,000 = $274,000 DO you agree with that?

OpenStudy (pphalke):

yes I do agree

OpenStudy (wolf1728):

Here's the formula I use for mortgages:

OpenStudy (pphalke):

ok

OpenStudy (wolf1728):

The rate would be 5.25 / 1,200 = 0.004375 principal = 274,000 months = 20 years * 12 = 240 Now, we put those numbers into the formula.

OpenStudy (wolf1728):

Monthly Payment = {0.004375 + 0.004375 / [(1.004375)^240 -1]} * 274,000 Now we calculate that.

OpenStudy (pphalke):

how do I calculate that I get 1668.88

OpenStudy (wolf1728):

No I'll show you

OpenStudy (wolf1728):

Payment = [0.004375 + {0.004375 / [(1.004375)^240 -1]}] * 274,000 Payment = {0.004375 + [0.004375 / 1.8511140206]} * 274,000 Payment = {0.004375 + 0.0023634417} * 274,000 Payment = 1846.3330157857 Payment = 1846.33 (rounded)

OpenStudy (wolf1728):

And to double check, here's a loan payment calculator. http://www.1728.org/calcloan.htm